How much pure alcohol must I add to 1 litter of 15% alcohol to produce a 35% alcohol solution?
Answer by lwsshak3(11628)

How much pure alcohol must I add to 1 litter of 15% alcohol to produce a 35% alcohol solution?
**
let x= liters of pure alcohol to be added
15%*1+100%*x=35%(x+1)
.15+x=.35x+.35
.65x=.20
x=20/65=4/13
ans:
liters of pure alcohol to be added=4/13